sketch and identify the curve defined by the parametric equations
(I). y=t+1 (II). x=t^2-2t (III). -4_< t _<6
1
Expert's answer
2020-07-09T19:48:09-0400

"y=t+1,x=t^2-2t, -4\le t \le6"


"t=y-1,\nx=(y-1)^2-2(y-1)"


"x=y^2-4y+3, -3\le y\le7"


This curve is parabola with a vertex at a point "(-1,2)"